# Quantum reflection time and the Goos-Hänchen effect

### v1

@@ I Introduction
In this work, we revisit the quantum Goos-Hänchen effect in order to establish whether determination of the GH shift does amount to measuring the time spent by the particle in the scattering potential. We will answer the question by the negative, and proceed to determine what is being measured, and to what accuracy. The paper is organised as follows. In Section II, we will rewrite the reflected wave packet (WP) in terms of the freely propagating counterpart, and a number of its mirror images. In section III we exploit the analogy with a von Neumann quantum measurement performed on a pre- and post-selected quantum system, and identify the measured quantity. In Sect. IV we show that in the classical limit one does determine the duration spent by the classical trajectory in the scattering potential. In Sect. V we attempt to extend the idea to the fully quantum case, and fail. In Sect. VI obtain the reflection “phase time” and discuss its properties. Sects. VII and VIII offer two simple examples of what has been said so far. In Sect. IX we abandon the attempts to define a meaningful “reflection time”, and replace instead a soft potential barrier by a superposition of virtual hard walls. In Sect. X, we use this representation to reconstruct the reflected beam, and evaluate the GH shift in terms of the spatial shifts experienced by the reflected particle. Section XI contains our conclusions.
@@ II Reflection in the momentum and coordinate representations
Figure 1: a) A wave packet with a mean momentum $p$, incident on a potential barrier $V(x)$, is partly reflected. Its mean energy is lower than the barrier height, and $x_{<}$ denotes the classical turning point $V(x_{<})=E(p)$. b) Freely propagating WP and its reflection is a mirror placed at $x=x^{\prime}/2$. c) The reflected WP shown in (a) is a superposition of the WPs reflected by a variety of infinite walls.
@@
$\Psi(x,t)=\Psi_{0}(x,t)+\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad$ (2)
@@
is the freely evolving initial state, and $R(k)$ is the reflection amplitude for a particle with momentum $k$, incident on the potential from the left. The function $R(k)$ is known to be analytic in the complex $k$-plane, with poles in its lower half, and also on the positive imaginary axis, if $V(x)$ supports bound states BZP . Note that Eq.(2) is valid even when if the initial state (1) contains negative momenta, in which case one uses $R(-|k|)=R^{*}(|k|)$ (see Appendix A).
@@
The evolution of the wave function in the left half-space can now be visualised as follows. At first, the WP localised far to the left, moves freely towards the scatterer, and the wave function is given by the first term in Eq.(2), $\Psi(x,t)\approx\Psi_{0}(x,t)$. For a while its mirror images $\Psi_{0}(x^{\prime}-x,t)$, moving in the opposite direction, remain “off screen” in the right half-space $x>0$. As the free WP crosses the origin, and disappears from the screen, its mirror images enter the region $x<0$, replacing there $\Psi_{0}(x,t)$. The scattering is finished when the left half-space contains only the reflected (R) part of the wave function
@@ III Analogy with a quantum measurement
Let the initial state of a particle of mass $m$ be a Gaussian WP with a mean momentum $p=mv>0$, and coordinate width $\Delta x$, placed a $x=x_{0}<0$. Throughout the paper we will neglect the WP’s spreading, so the free state in Eq.(3) propagates without distortion, (see Appendix B), and we have
@@
$\Psi_{R}(x,t)\approx\exp[-ip(x+x_{0})-iE(p)t]\times\quad\quad\quad\quad\quad$ (11)
@@
The usual dilemma is now evident. In the von Neumann example of Appendix D, interaction with the pointer alters the likelihood of the post-selection. Similarly, the reflection probability $P_{R}(t)\equiv\int dx|\Psi_{R}(x,t)|^{2}$ depends on the WP’s width $\Delta x$. An accurate measurement, $\Delta x\to 0$, destroys reflection, $P_{R}(t)\to 0$, since most of the WP broad in the momentum space, and is now transmitted. Conversely, the choice of a nearly monochromatic WP, $\Delta x\to\infty$, preserves $P_{R}(t)\to|R(p)|^{2}$ but fails to distinguish between different shifts $x^{\prime}$. A similar claim that, in a double slit experiment, one cannot both know the slit chosen by the particle and retain the interference pattern on the screen is often used to illustrate the Uncertainty Principle FeynL .
@@
Consider the semiclassical limit of Eqs.(11) in the case of the barrier shown in Fig.1a. For energies less than the barrier height the reflection amplitude is given by Land
@@
${\langle}x{\rangle}\equiv\frac{\int_{-\infty}^{0}dx\,x\,|\Psi_{R}(x,t)|^{2}}{\int_{-\infty}^{0}dx\,\Psi_{R}(x,t)|^{2}}$ (18)
@@ V Quantum “reflection time”
where $\phi_{R}(p)$ is the phase of the reflection coefficient, $R(p)=|R(p)|\exp[i\phi_{R}(p)]$. The phase time (23) is often thought to represent the duration a particle with momentum $p$ spends in the reflecting potential. [In quantum measurements, the complex quantity in the brackets is known as “weak value of $x^{\prime}$W1 , W2 . It can be shown that, as in the case of transmission W3 , its imaginary part determines the small change in the reflected particle’s mean momentum, which occurs because higher momenta are transmitted more easily, and are less likely to be reflected (see also GH4 ).]
@@
For a particle scattered by a zero-range potential $V(x)=\Omega\delta(x)$, where $\delta(x)$ is the Dirac delta, the reflection amplitude is well-known to be
@@
It is readily seen that $\int_{0}^{\infty}dx^{\prime}\xi(x^{\prime})=-1$ and, as the height of the step increases, $\xi(x^{\prime}){\xrightarrow[V\to\infty]{}}-\delta(x^{\prime})$, and the reflection amplitude tends to that of a hard wall,
@@
$\tau_{ph}(p)=\begin{cases}\frac{2m}{p\sqrt{2mV-p^{2}}}&\text{ for}\quad 0\leq p\leq\sqrt{2mV}\\ 0&\text{for}\quad p>\sqrt{2mV},\end{cases}$ (33)
@@ X Quantum Goos-Hänchen effect
$\Psi^{\infty}_{R}(x,y,t|a)=\Psi_{0}(2a-x,y,t).$ (40)
Let $\Psi_{0}(x,y,t)$ be a transverse Gaussian beam (or a WP very broad in the direction of propagation), making an angle $\theta$ with the normal to the surface $x=0$ (see Fig. 5a). In the coordinate system $(x_{1},y_{2})$, obtained by rotating the axes $(x,y)$ counter clockwise by $\theta$, we have
@@
The quantity of interest is the Goos-Hänchen shift $\delta y_{GH}$, i.e., the distance by which the reflected beam is displaced relative to the same beam reflected by a hard wall placed at $x=0$ (see Fig.2b). We note that superposition of the Gaussians in Eq.(41) does not have to be a Gaussian itself. However, one can always calculate the centre of the beam, the way it was done in Eq.(19)
@@
$S_{GH}=\frac{{\langle}y_{2}{\rangle}}{\cos\theta}.$ (45)
@@
$S_{GH}{\xrightarrow[\Delta y/\sin\theta\to 0]{}}\tan(\theta)\frac{\int_{0}^{\infty}dx^{\prime}x^{\prime}|\xi(x^{\prime})|^{2}}{\int_{0}^{\infty}dx^{\prime}|\xi(x^{\prime})|^{2}}$ (46)
@@
$S_{GH}{\xrightarrow[\Delta y/\sin\theta\to\infty]{}}\tan(\theta)\text{Re}\left[\frac{\int_{0}^{\infty}dx^{\prime}x^{\prime}\exp(ip_{x}x^{\prime})\xi(x^{\prime})}{\int_{0}^{\infty}dx^{\prime}\exp(ip_{x}x^{\prime})\xi(x^{\prime})}\right]\quad$ (47)
@@
Finally, in the paraxial approximation of Appendix G, effects of beam’s broadening can be included by replacing $G_{0}(y)$ and $\tilde{G}_{0}(q)$ In Eqs.(41) and (48), by their coordinate-dependents versions (91) and (93). This approach is also useful beyond the paraxial approximation. For example, consider a doubly Gaussian wave packet, reflected by potential step (see section VIII) in the regime of total reflection (i.e when all energies in the wave packet are much lower than the step height). In this case, both the longitudinal delay and the lateral displacement can be directly related to the average depth at which the particle is reflected,
@@ XI Conclusions
Above we attempted to give a most detailed description of the quantum Goos-Hänchen effect that an elementary theory can offer. For a classical particle with a known momentum launched at oblique angle at a repulsive potential, the distance $S_{GH}$ between the exit and entry point at the surface, where the potential vanishes, can be found by multiplying its lateral velocity $v_{II}$ by the duration $\tau$ its trajectory spends in the potential: $S_{GH}=v_{II}\tau$.
A straightforward extension of this analysis to the quantum case is, however, not possible. In one dimension, the quantum state of a particle with a known momentum $p=mv$ is a plane wave, and the moments when it enters and leaves the potential region simply cannot be defined. One notes, however, that the reflected state is a superposition (13) of multiple plane waves with phase shits corresponding to displacements, or shifts, $x^{\prime}$, induced by the potential. One notes also that in experiment where one prepares the particle in a wave packet (WP) state with a mean momentum $p$, and then observes it at $x$, the shifts are restricted to the range, determined by the WP’s coordinate width $\Delta x$. In the classical limit of this “quantum measurement”, the unique displacement $x^{\prime}_{s}$, with which the reflected particle leaves the potential, corresponds to the stationary region where rapid oscillations of the amplitude distribution $\xi(x^{\prime})$ slow down. This classical value allows one to deduce the duration $\tau$ that the particle’s trajectory spends in the potential as $x^{\prime}_{s}/v$.
There are two good reasons why one cannot continue to use the time argument beyond the classical limit, e.g., by converting each spatial shift into a duration using $\tau(x^{\prime})=x^{\prime}/v$. One is that some of these “durations” will have to be negative in the case of a potential well, or, indeed, of any potential supporting bound states. The idea that a particle which which makes a U-turn before entering the potential spends in the scatterer a “negative duration” does not help explain the phenomenon in classical-like terms. [We are aware of the recently reported “experimental evidence of negative times” St , and remain sceptical.]. The notion that a particle’s state can acquire a negative phase shift can, however, be made without causing a further controversy.
@@
For example, the hard wall picture helps one to visualise what happens in the quantum Goos-Hänchen case GH1 . The outgoing Gaussian beam is composed of the beams reflected by each fictitious wall (see Fig. 5.a). The cross section of the beam’s wave function is a superposition of Gaussians (we neglect the lateral spreading of the wavepacket during propagation) similar to one would obtain in a von Neumann quantum measurement. As the width of the incident beam increases (the measurement becomes less accurate) the Goos-Hänchen shift changes from Eq.(46) to Eq.(61), involving the “strong” and the “weak” values of the coordinate shift $x^{\prime}$.
@@ Appendix A. Scattering by a one-dimensional barrier
${\langle}{x}|\Phi_{l}(k){\rangle}=\frac{1}{\sqrt{2\pi}}\left[e^{ikx}+R_{l}(k)e^{-ikx}\right],\quad x<0\quad$ (52)
$=\frac{1}{\sqrt{2\pi}}T_{l}(k)e^{ikx},\quad x>d\quad$
@@
$\int_{0}^{\infty}dk\left[|\Phi_{l}(k){\rangle}{\langle}\Phi_{l}(k)|+|\Phi_{r}(k){\rangle}{\langle}\Phi_{r}(k)|\right]=1.$ (53)
@@
$T_{l}(k)=T_{r}(k)=T(k),\quad R_{l}(k)T^{*}(k)=-R^{*}_{r}(k)T(k),\quad\quad$ (54)
$T(-k^{*})=T(k)^{*},\quad R_{l.r}(-k^{*})=R^{*}_{l.r}(k).\quad\quad\quad\quad\quad\quad\quad$
@@
$\int_{0}^{\infty}dk{\langle}\Phi_{l}(k)|\Psi(0){\rangle}e^{-iE(k)t}{\langle}x|\Phi_{l}(k){\rangle},$
@@
${\langle}\Phi_{l}(k)|\Psi(0){\rangle}=(2\pi)^{-1}\int_{-\infty}^{\infty}dk^{\prime}\int_{-\infty}^{\infty}dx^{\prime}\quad\quad\quad$ (58)
$\left[e^{-ikx^{\prime}}+R_{l}^{*}(k)e^{ikx^{\prime}}\right]e^{ik^{\prime}x^{\prime}}{\langle}k^{\prime}|\Psi(0){\rangle}=\quad\quad\quad\quad\quad$
${\langle}k|\Psi(0){\rangle}+R_{l}^{*}(k){\langle}-k|\Psi(0){\rangle},\quad\quad\quad\quad\quad\quad\quad\quad$
${\langle}\Phi_{r}(k)|\Psi(0){\rangle}=(2\pi)^{-1}\int_{-\infty}^{\infty}dk^{\prime}\int_{-\infty}^{\infty}dx^{\prime}\quad\quad\quad$
$T^{*}(k)e^{ikx^{\prime}}e^{ik^{\prime}x^{\prime}}{\langle}k^{\prime}|\Psi(0){\rangle}=T^{*}(k){\langle}-k|\Psi(0){\rangle}.\$
[…]
$\Psi(x,t)=\frac{1}{\sqrt{2\pi}}\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad$ (59)
$\{\int_{0}^{\infty}dk[{\langle}k|\Psi(0){\rangle}+R_{l}^{*}(k){\langle}-k|\Psi(0){\rangle}]T(k)e^{ikx-iE(k)t}$
$+\int_{0}^{\infty}dkT^{*}(k){\langle}-k|\Psi(0){\rangle}[e^{-ikx}+R_{r}(k)e^{ikx}]e^{-iE(k)t}\},$

### v2

@@ I Introduction
In this work, we revisit the quantum Goos-Hänchen effect in order to establish whether determination of the GH shift does amount to measuring the time spent by the particle in the scattering potential. We will answer the question by the negative, and proceed to determine what is being measured, and to what accuracy. The paper is organised as follows. In Section II, we will rewrite the reflected wave packet (WP) in terms of the freely propagating counterpart, and a number of its mirror images. In section III we exploit the analogy with a von Neumann quantum measurement performed on a pre- and post-selected quantum system, and identify the measured quantity. In Sect. IV we show that in the classical limit one does determine the duration spent by the classical trajectory in the scattering potential. In Sect. V we attempt to extend the idea to the fully quantum case, and fail. In Sect. VI we obtain the reflection “phase time” and discuss its properties. Sects. VII and VIII offer two simple examples of what has been said so far. In Sect. IX we abandon the attempts to define a meaningful “reflection time”, and replace instead a soft potential barrier by a superposition of virtual hard walls. In Sect. X, we use this representation to reconstruct the reflected beam, and evaluate the GH shift in terms of the spatial shifts experienced by the reflected particle. Section XI contains our conclusions.
@@ II Reflection in the momentum and coordinate representations
Figure 1: a) A wave packet with mean momentum $p$, incident on a potential barrier $V(x)$, is partly reflected. Its mean energy is lower than the barrier height, and $x_{<}$ denotes the classical turning point $V(x_{<})=E(p)$. b) Freely propagating WP and its reflection is a mirror placed at $x=x^{\prime}/2$. c) The reflected WP shown in (a) is a superposition of the WPs reflected by a variety of infinite walls.
@@
$\Psi(x,t)=\Psi_{0}(x,t)+$ (2)
@@
is the freely evolving initial state, and $R(k)$ is the reflection amplitude for a particle with momentum $k$, incident on the potential from the left. The function $R(k)$ is known to be analytic in the complex $k$-plane, with poles in its lower half, and also on the positive imaginary axis, if $V(x)$ supports bound states BZP . Note that Eq.(2) is valid even when the initial state (1) contains negative momenta, in which case one uses $R(-|k|)=R^{*}(|k|)$ (see Appendix A).
@@
The evolution of the wave function in the left half-space can now be visualised as follows. At first, the WP localised far to the left, moves freely towards the scatterer, and the wave function is given by the first term in Eq.(2), $\Psi(x,t)\approx\Psi_{0}(x,t)$. For a while its mirror images $\Psi_{0}(x^{\prime}-x,t)$, moving in the opposite direction, remain “off screen” in the right half-space $x>0$. As the free WP crosses the origin, and disappears from the screen, its mirror images enter the region $x<0$, replacing there $\Psi_{0}(x,t)$. Scattering is finished when the left half-space contains only the reflected (R) part of the wave function,
@@ III Analogy with a quantum measurement
Let the initial state of a particle of mass $m$ be a Gaussian WP with a mean momentum $p=mv>0$, and coordinate width $\Delta x$, placed at $x=x_{0}<0$. Throughout the paper we will neglect the WP’s spreading, so the free state in Eq.(3) propagates without distortion, (see Appendix B), and we have
@@
$\Psi_{R}(x,t)$ $\approx\exp[-ip(x+x_{0})-iE(p)t]\times$ (11)
@@
The usual dilemma is now evident. In the von Neumann example of Appendix D, interaction with the pointer alters the likelihood of the post-selection. Similarly, the reflection probability $P_{R}(t)\equiv\int dx|\Psi_{R}(x,t)|^{2}$ depends on the WP’s width $\Delta x$. An accurate measurement, $\Delta x\to 0$, destroys reflection, $P_{R}(t)\to 0$, since most of the WP, broad in the momentum space, and is now transmitted. Conversely, the choice of a nearly monochromatic WP, $\Delta x\to\infty$, preserves $P_{R}(t)\to|R(p)|^{2}$ but fails to distinguish between different shifts $x^{\prime}$. A similar claim that, in a double slit experiment, one cannot both know the slit chosen by the particle and retain the interference pattern on the screen is often used to illustrate the Uncertainty Principle FeynL .
@@
Consider the semiclassical limit of Eq.(11) in the case of the barrier shown in Fig.1a. For energies less than the barrier height the reflection amplitude is given by Land
@@
${\langle}x{\rangle}\equiv\frac{\int_{-\infty}^{0}dx\,x\,|\Psi_{R}(x,t)|^{2}}{\int_{-\infty}^{0}dx|\Psi_{R}(x,t)|^{2}}$ (18)
@@ V Quantum “reflection time”
where $\phi_{R}(p)$ is the phase of the reflection coefficient, $R(p)=|R(p)|\exp[i\phi_{R}(p)]$. The phase time (23) is often thought to represent the duration a particle with momentum $p$ spends in the reflecting potential. [In quantum measurements, the complex quantity in the brackets is known as “weak value of $x^{\prime}$W1 , W2 . It can be shown that, as in the case of transmission DS3 , its imaginary part determines the small change in the reflected particle’s mean momentum, which occurs because higher momenta are transmitted more easily, and are less likely to be reflected (see also GH4 ).]
@@
For a particle scattered by a zero-range potential $V(x)=\Omega\delta(x)$, where $\delta(x)$ is the Dirac delta, the reflection amplitude is well known to be
@@
It is readily seen that $\int_{0}^{\infty}dx^{\prime}\xi(x^{\prime})=-1$ and, as the height of the step increases, $\xi(x^{\prime}){\xrightarrow[V\to\infty]{}}{-\delta(x^{\prime})}$, the reflection amplitude tends to that of a hard wall,
@@
$\tau_{ph}(p)=\begin{cases}\frac{2m}{p\sqrt{2mV-p^{2}}}&\text{ for}\quad 0\leq p\leq\sqrt{2mV}\\ 0&\text{for}\quad p>\sqrt{2mV},\quad\end{cases}$ (33)
@@ X Quantum Goos-Hänchen effect
$\Psi^{\infty}_{R}(x,y,t|a)=-\Psi_{0}(2a-x,y,t).$ (40)
Let $\Psi_{0}(x,y,t)$ be a transverse Gaussian beam (or a WP very broad in the direction of propagation), making an angle $\theta$ with the normal to the surface $x=0$ (see Fig. 5a). In the coordinate system $(x_{1},y_{1})$, obtained by rotating the axes $(x,y)$ counter clockwise by $\theta$, we have
@@
The quantity of interest is the Goos-Hänchen shift $\delta y_{\text{GH}}$, i.e., the distance by which the reflected beam is displaced relative to the same beam reflected by a hard wall placed at $x=0$ (see Fig.5b). We note that superposition of the Gaussians in Eq.(43) does not have to be a Gaussian itself. However, one can always calculate the centre of the beam, the way it was done in Eq.(19)
@@
$S_{\text{GH}}=\frac{{\langle}y_{2}{\rangle}}{\cos(\theta)}.$ (45)
@@
$S_{\text{GH}}{\xrightarrow[\Delta y/\sin\theta\to 0]{}}\tan(\theta)\frac{\int_{0}^{\infty}dx^{\prime}x^{\prime}|\xi(x^{\prime})|^{2}}{\int_{0}^{\infty}dx^{\prime}|\xi(x^{\prime})|^{2}}$ (46)
@@
$S_{\text{GH}}{\xrightarrow[\Delta y/\sin\theta\to\infty]{}}\tan(\theta)\text{Re}\left[\frac{\int_{0}^{\infty}dx^{\prime}x^{\prime}\exp(ip_{x}x^{\prime})\xi(x^{\prime})}{\int_{0}^{\infty}dx^{\prime}\exp(ip_{x}x^{\prime})\xi(x^{\prime})}\right]\quad\quad$ (47)
@@
Finally, in the paraxial approximation of Appendix G, effects of beam’s broadening can be included by replacing $G_{0}(y)$ and $\tilde{G}_{0}(q)$ in Eqs.(41) and (48), by their coordinate-dependents versions (93) and (95). This approach is also useful beyond the paraxial approximation. For example, consider a doubly Gaussian wave packet, reflected by potential step (see Section VIII) in the regime of total reflection (i.e., when all energies in the wave packet are much lower than the step height). In this case, both the longitudinal delay and the lateral displacement can be directly related to the average depth at which the particle is reflected,
@@ XI Conclusions
Above we attempted to give a most detailed description of the quantum Goos-Hänchen effect that an elementary theory can offer. For a classical particle with a known momentum launched at oblique angle at a repulsive potential, the distance $S_{\text{GH}}$ between the exit and entry point at the surface, where the potential vanishes, can be found by multiplying its lateral velocity $v_{\parallel}$ by the duration $\tau$ its trajectory spends in the potential: $S_{\text{GH}}=v_{\parallel}\tau$.
A straightforward extension of this analysis to the quantum case is, however, not possible. In one dimension, the quantum state of a particle with a known momentum $p=mv$ is a plane wave, and the moments when it enters and leaves the potential region simply cannot be defined. One notes, however, that the reflected state is a superposition (13) of multiple plane waves with phases corresponding to spatial displacements, or shifts, $x^{\prime}$, induced by the potential. One notes also that in experiment where one prepares the particle in a wave packet state with a mean momentum $p$, and then observes it at $x$, the shifts are restricted to the range, determined by the WP’s coordinate width $\Delta x$. In the classical limit of this “quantum measurement”, the unique displacement $x^{\prime}_{s}$, with which the reflected particle leaves the potential, corresponds to the stationary region where rapid oscillations of the amplitude distribution $\xi(x^{\prime})$ slow down. This classical value allows one to deduce the duration $\tau$ that the particle’s trajectory spends in the potential as $x^{\prime}_{s}/v$.
There are two good reasons why one cannot continue to use the time argument beyond the classical limit, e.g., by converting each spatial shift into a duration using $\tau(x^{\prime})=x^{\prime}/v$. One is that some of these “durations” will have to be negative in the case of a potential well, or, indeed, of any potential supporting bound states. The idea that a particle which makes a U-turn before entering the potential spends in the scatterer a “negative duration” does not help explain the phenomenon in classical-like terms. [We are aware of the recently reported “experimental evidence of negative times” St , and remain sceptical.] The notion that a particle’s state can acquire a negative phase shift can, however, be made without causing a further controversy.
@@
For example, the hard wall picture helps one to visualise what happens in the quantum Goos-Hänchen case GH1 . The outgoing Gaussian beam is composed of the beams reflected by each fictitious wall (see Fig.5a). The cross section of the beam’s wave function is a superposition of Gaussians (we neglect the lateral spreading of the wavepacket during propagation) similar to one would obtain in a von Neumann quantum measurement. As the width of the incident beam increases (the measurement becomes less accurate) the Goos-Hänchen shift changes from Eq.(46) to Eq.(47), involving the “strong” and the “weak” values of the coordinate shift $x^{\prime}$.
@@ Appendix A. Scattering by a one-dimensional barrier
${\langle}{x}|\Phi_{l}(k){\rangle}=\frac{1}{\sqrt{2\pi}}\left[e^{ikx}+R_{l}(k)e^{-ikx}\right],\quad$ $x<0$ (52)
$=\frac{1}{\sqrt{2\pi}}T_{l}(k)e^{ikx},\quad$ $x>d$
@@
$\int_{0}^{\infty}dk\Big[|\Phi_{l}(k){\rangle}{\langle}\Phi_{l}(k)|+|\Phi_{r}(k){\rangle}{\langle}\Phi_{r}(k)|\Big]=1.$ (53)
@@
$T_{l}(k)=T_{r}(k)=T(k),\quad$ $R_{l}(k)T^{*}(k)=-R^{*}_{r}(k)T(k),$ (54)
$T(-k^{*})=T(k)^{*},\quad$ $R_{l,r}(-k^{*})=R^{*}_{l,r}(k).$
@@
$\int_{0}^{\infty}dk{\langle}\Phi_{l}(k)|\Psi(0){\rangle}e^{-iE(k)t}{\langle}x|\Phi_{l}(k){\rangle},$ (57)
@@
${\langle}\Phi_{l}(k)|\Psi(0){\rangle}=(2\pi)^{-1}\int_{-\infty}^{\infty}dk^{\prime}\int_{-\infty}^{\infty}dx^{\prime}$ (59)
$\left[e^{-ikx^{\prime}}+R_{l}^{*}(k)e^{ikx^{\prime}}\right]e^{ik^{\prime}x^{\prime}}{\langle}k^{\prime}|\Psi(0){\rangle}=$
${\langle}k|\Psi(0){\rangle}+R_{l}^{*}(k){\langle}-k|\Psi(0){\rangle},$
${\langle}\Phi_{r}(k)|\Psi(0){\rangle}=(2\pi)^{-1}\int_{-\infty}^{\infty}dk^{\prime}\int_{-\infty}^{\infty}dx^{\prime}$
$T^{*}(k)e^{ikx^{\prime}}e^{ik^{\prime}x^{\prime}}{\langle}k^{\prime}|\Psi(0){\rangle}=T^{*}(k){\langle}-k|\Psi(0){\rangle}.$
Inserting (59) into (56) for $x>d$ yields
$\Psi$ $(x,t)=\frac{1}{\sqrt{2\pi}}\times$ (60)
$\Big\{$ $\int_{0}^{\infty}dk[{\langle}k|\Psi(0){\rangle}+R_{l}^{*}(k){\langle}-k|\Psi(0){\rangle}]T(k)e^{ikx-iE(k)t}$
$+$ $\int_{0}^{\infty}dkT^{*}(k){\langle}-k|\Psi(0){\rangle}[e^{-ikx}+R_{r}(k)e^{ikx}]e^{-iE(k)t}\Big\},$